# Reinforcing Clinical Decision Support through Multi-Agent Systems and Ethical AI Governance

### v2

@@ Abstract
In the age of data-driven medicine, it is paramount to include explainable and ethically managed artificial intelligence in explaining clinical decision support systems to achieve trustworthy and effective patient care. The focus of this paper is on a new architecture of a multi-agent system for clinical decision support that uses modular agents to analyze laboratory results, vital signs, and the clinical context and then integrates these results to drive predictions and validate outcomes. We describe our implementation with the eICU database to run lab-analysis-specific agents, vitals-only interpreters, and contextual reasoners and then run the prediction module and a validation agent. Everything is a transparent implementation of business logic, influenced by the principles of ethical AI governance such as Autonomy, Fairness, and Accountability. It provides visible results that this agent-based framework not only improves on interpretability and accuracy but also on reinforcing trust in AI-assisted decisions in an intensive care setting.
@@
Clinical Decision Support, Mortality Prediction in the ICU, Deep Learning, Transparency, Large Language Model, AI Agent
@@ I Introduction
Artificial intelligence (AI) has steadily made its way into many industries, from image recognition and supply chain logistics [1, 2, 3], writing assist [4] to the field of healthcare [5, 6]. Within medicine, it’s proving valuable for sharpening diagnostic precision, supporting treatment planning, and helping clinicians keep a closer eye on patients [7]. A growing body of work has focused on using AI to interpret complex medical visuals like surgical footage [8], X-rays [9], computed tomography (CT) scans [10], and magnetic resonance imaging (MRI) scans [11], making interpretation faster and more consistent. But it doesn’t stop with images. AI is also being used to make sense of physiological signals, including electroencephalography (EEG) [12, 13], electrocardiography (ECG) [14], and data from wearable sensors [15, 16]. These efforts are opening up new possibilities across neurology, psychiatry, and continuous patient monitoring [17]. Altogether, these advancements point to a future where AI supports both visual and signal-based insights, forming the backbone of smarter clinical decision-making tools.
Clinical decision support systems (CDSS) have become a vital part of today’s healthcare settings, offering insights drawn from electronic health records (EHRs) and real-time monitoring tools. Yet, many of the traditional AI models used in these systems fall short when it comes to flexibility, transparency, and oversight key qualities, especially critical in high-risk settings like intensive care units (ICUs). To address these limitations, we introduce a modular multi-agent system (MAS) designed to reflect how clinical teams make decisions, with a built-in emphasis on ethical AI to uphold both explainability and accountability.
Building on progress in agent-based frameworks and the coordinated use of large language models (LLMs), our system breaks down the decision-making pipeline into focused, collaborative agents. Each one is responsible for a different aspect of ICU assessment: from interpreting lab results and tracking vital signs to making context-sensitive judgments based on a patient’s history or co-existing conditions. These individual agents pass their findings to a central integration agent that brings everything together, enabling more comprehensive predictions and cross-validated outcomes. This structure mimics how doctors gathering evidence from various sources, weighing context, and forming a unified clinical picture.
By structuring the system around modular agents and grounding it in ethical oversight, we improve not just how interpretable and scalable the model is, but also how it upholds fairness and accountability throughout the clinical decision-making process. To test the framework, we draw on the eICU Collaborative Research Database [18], showing that our method can deliver well-organized predictions, shed light on key prognostic indicators, and build greater trust in AI-supported medical judgments.
@@ II-A Applications of Clinical Decision Support Systems in Intensive Care Settings
Clinical decision support systems (CDSS) have come a long way, especially in ICU environments where every second counts. Earlier systems typically leaned on rule-based logic or statistical methods to generate recommendations [19, 20]. More recent developments have looked to clinical practice guidelines (CPGs) as a way to enrich LLMs, boosting their ability to offer context-aware treatment advice. Research suggests that LLMs enhanced with CPGs outperform traditional models in delivering more accurate clinical suggestions [19].
Meanwhile, multi-agent approaches to CDSS have gained traction as well. One notable design introduced a case-based reasoning (CBR) framework structured around agents that handle user interaction, task execution, and domain knowledge [20]. Combining MAS with CBR has been shown to improve how efficiently the system learns and adapts to the unique traits of each patient.
@@
The electronic ICU relies on a telemedicine framework to monitor high-risk patients and deliver critical care, even when onsite specialists are not available [22]. For instance, Philips has developed an eICU system that utilizes the eCareManager platform to bring expert ICU support directly to the patient’s bedside. By linking hospital networks and supplying real-time clinical feedback, this approach narrows the gap between off-site experts and immediate patient needs [23].
In practical terms, the rollout of eICU systems has produced measurable improvements in patient outcomes. At Baptist Health South Florida, the introduction of the eICU model was linked to a 23% decrease in ICU mortality and a reduction in the average length of stay by up to 25% [23]. These results not only demonstrate how telemedicine can streamline critical care delivery but also highlight the potential for such systems to transform clinical practice through enhanced real-time decision-making and optimized resource use.
@@ II-C LLM-Based Agents in Healthcare
Large language models (LLMs) have recently become more common in healthcare. These models now assist in multiple areas, including virtual assistants, individualized health education, symptom checking, and mental health support tools [24]. By improving patient interactions and simplifying administrative tasks, LLM-based systems are beginning to influence how healthcare is provided.
One significant example is MDAgents, a multi-agent system using LLMs to manage complex medical decisions [25]. Its design replicates the teamwork observed in actual healthcare environments, enabling effective communication among its agents. Testing has shown that MDAgents performs better than earlier models in various evaluations.
A recent review explored the use of LLM-based agents in medical contexts [26]. The review covered technical foundations, practical applications, and existing challenges. It emphasized components such as planning techniques, reasoning strategies, integration of external tools, and agent architecture. These systems are now employed for tasks like clinical decision support, automatic patient documentation, simulation training, and workflow optimization.
Researchers are now further developing LLM-based agents into multi-agent systems (MAS), where multiple agents interact in a decentralized and collaborative way. This change allows for systems that are more organized and flexible, offering new ways to manage challenging healthcare situations, such as emergency response coordination and personalized patient treatment.
@@ II-D Multi-Agent Systems in Healthcare
Multi-agent systems (MAS) are gaining attention as a promising way to tackle complex challenges in healthcare. One example applies MAS to pre-hospital emergency response, where agents—such as EMS dispatch centers, ambulances, traffic nodes, and medical providers—collaborate within a distributed decision-making setup [27].
The idea of multi-stage AI agents builds on this by organizing intelligent agents into layers, each handling different parts of perception and reasoning. Many of these layers are now powered by LLMs, allowing for more structured and scalable workflows [28]. This layered setup has shown particular promise in areas like personalized care and remote health services.
Real-world implementations are emerging as well. The LLM-medical-agent framework, for instance, demonstrates how MAS can be applied to modular analysis of healthcare data in practical settings [29].
@@ II-E Ethical Governance in Healthcare AI
As AI continues to find its place in healthcare, the need for ethical oversight and transparent reasoning becomes more urgent. Explainable AI (XAI) plays a key role here by making machine-generated decisions easier for humans to understand—helping both clinicians and patients build confidence in AI-supported care [30]. By shedding light on the logic behind predictions, XAI tackles the long-standing “black-box” issue in conventional AI systems.
Growing concerns around the safety of LLM-based agents have prompted the development of frameworks like GuardAgent [31], which embed policy guardrails to ensure compliance with safety and privacy standards—an especially important safeguard in clinical environments [32].
@@
To that end, the World Health Organization (WHO) has outlined ethical guidelines for AI in healthcare, highlighting principles such as human autonomy, wellbeing, and system transparency [34]. Especially in high-risk areas like intensive care units, addressing these governance challenges is key to the responsible deployment of AI [35].
@@ II-F Motivation and Research Gap
While clinical decision support systems, LLM-powered agents and multi-agent frameworks [36], have made notable strides in other fields, there’s still a considerable gap when it comes to integrating these technologies into healthcare, especially for real-world ICU settings under clear ethical oversight. Many current solutions fall short in modularity, lack transparency, or aren’t built with the kind of structured, inter-agent communication needed to reflect the fast-paced, interdisciplinary nature of intensive care.
Most existing research tends to focus on isolated tasks—like interpreting lab results, monitoring vital signs, or reasoning based on medical history—but few bring these components together into a unified system that reflects how clinicians actually work as a team. This fragmentation underscores the need for a new approach.
In response, we propose a novel agent-based architecture that breaks down the clinical reasoning process into specialized, collaborative agents—each designed to handle a distinct aspect of care while maintaining accountability and interpretability throughout. By embedding ethical AI principles directly into each stage of the pipeline and validating our design using the eICU database, we aim to bridge both the technical and ethical gaps in deploying trustworthy AI for high-stakes decision-making in critical care.
Figure 1: Illustration of the Multi-Agent Framework Design. The system consists of a set of specialized agents, each responsible for processing a specific type of clinical data. The Context Analysis Agent handles unstructured inputs like clinical notes, while the Vitals Analysis Agent focuses on real-time physiological signals, and the Lab Analysis Agent interprets laboratory test results. These distinct streams of information are brought together by the Integration Agent, which fuses multimodal features into a unified representation. Based on this, the Prediction Agent carries out key forecasting tasks—such as predicting ICU mortality or estimating length of stay. To support interpretability, the Transparency Agent generates human-readable explanations of model outputs. Finally, the Validation Agent oversees performance assessment by comparing predictions against ground truth data.
@@ III-A Dataset and Preprocessing
This study used the eICU Collaborative Research Database v2.0 [21], which compiles anonymized ICU records from over 200,000 patient admissions in diverse U.S. hospitals. The database encompasses structured details (e.g., vital signs and lab results) and unstructured clinical notes contributed by nurses and physicians, giving us a thorough view of patient care.
We concentrated on several key eICU files: patient.csv, lab.csv, vitalPeriodic.csv, note.csv, and medication.csv, along with APACHE-related data files [37] (apacheApsVar.csv and apachePatientResult.csv). To align each patient’s information, records were grouped according to patient-unit-stayid. If any essential data were missing—such as vital signs, lab values, or clinical notes—we removed those entries to maintain reliability.
We addressed the data gaps, ordered events based on their timestamps, and shortened lengthy text fields to meet language model guidelines. Then, we sampled 150 patients for the study: 76 mortality patients and 74 survived patients.
Next, we extracted specific features from each patient’s record set. We captured the ten most recent vital sign readings to reveal each patient’s current physiological state. We also selected the latest distinct lab biomarkers deemed clinically relevant. When dealing with unstructured clinical documentation, we included up to three notes for every patient, focusing largely on entries written by physicians and nurses.
Our analysis also tracked the top 20 medications and treatments, identifying them by frequency or uniqueness within the overall dataset. Finally, we incorporated APACHE scores and predictions as reference points to aid in validating and evaluating our modeling outcomes.
III-B Multi-Agent Framework Design
To emulate real-world ICU decision-making, we implemented a modular multi-agent architecture consisting of six discrete agents, each responsible for a semantically distinct task. The architecture is shown in Figure 1.
@@ III-B Multi-Agent Framework Design
  • Integration Agent: Aggregates all above agent outputs into a comprehensive, system-by-system clinical assessment. It prioritizes ICU risk factors related to mortality and length of stay (LOS).
  • Prediction Agent: Generates structured outcome predictions (mortality probability and ICU LOS) using integrated findings and APACHE variables. Outputs follow a strict template for automated parsing.
@@
To maximize information flow between agents, we implemented a shared memory architecture that allows any agent to access inputs and outputs from previous pipeline stages. This approach mitigates the risk of information loss between modules while maintaining the semantic separation of responsibilities. While this shared memory design has proven effective at improving predictive performance by ensuring that no critical clinical insights are lost during the analysis process, we recognize that the multi-agent architecture introduces additional complexity that can impact transparency. Our current implementation focuses on performance optimization, with ongoing work to enhance the explainability mechanisms across the agent communication pathways.
@@ III-D System Execution and DAG Orchestration
  • prediction feeds into validation.
@@
Figure 2: Comparison of performance between the Multi-agent and Single-agent frameworks across three evaluation metrics: Mortality Prediction Accuracy, Length of Stay (LOS) Mean Error, and LOS Median Error. Each model was executed 8 times, and the box plots represent the distribution of results over these runs. The Multi-agent framework shows slightly higher mean mortality accuracy with slightly more variance, while LOS-related errors are nearly identical between the two models. These results indicate comparable predictive performance, with minor differences in consistency and central tendency.
@@ III-F Ethical AI and Explainability
To ensure safety, fairness, and transparency, the system incorporates several governance mechanisms:
  • Explainability: All agent outputs follow enforced templates, allowing users to trace predictions back to specific findings.
  • Fairness: Demographic data (age, gender, ethnicity) are included in prompts to support bias analysis.
  • Autonomy Boundaries: Agents are not allowed to override downstream agents without structured justification.
  • Auditability: Logs, JSON outputs, and flow diagrams are archived per patient to support reproducibility.

### v3

# Enhancing Clinical Decision-Making: Integrating Multi-Agent Systems with Ethical AI Governance

@@ Abstract
Recent advances in the data-driven medicine approach, which integrates ethically managed and explainable artificial intelligence into clinical decision support systems (CDSS), are critical to ensure reliable and effective patient care. This paper focuses on comparing novel agent system designs that use modular agents to analyze laboratory results, vital signs, and clinical context, and to predict and validate results. We implement our agent system with the eICU database, including running lab analysis, vitals-only interpreters, and contextual reasoners agents first, then sharing the memory into the integration agent, prediction agent, transparency agent, and a validation agent. Our results suggest that the multi-agent system (MAS) performed better than the single-agent system (SAS) with mortality prediction accuracy (59%, 56%) and the mean error for length of stay (LOS)(4.37 days, 5.82 days), respectively. However, the transparency score for the SAS (86.21) is slightly better than the transparency score for MAS (85.5). Finally, this study suggests that our agent-based framework not only improves process transparency and prediction accuracy but also strengthens trustworthy AI-assisted decision support in an intensive care setting.
@@
Clinical Decision Support, Mortality Prediction in the ICU, Transparency, Length Of Stay, LLM Single Agent System, LLM Multi Agent System
@@ I Introduction
Artificial intelligence (AI) has been widely adopt into healthcare [1]. Within medicine, it’s proving valuable for sharpening diagnostic precision, supporting treatment planning, and helping clinicians take care of patients [2]. Recent work has focused on using AI to interpret complex medical visuals like surgical footage [3], computed tomography (CT) scans [4], and magnetic resonance imaging (MRI) scans [5], making interpretation faster and more consistent. These efforts provide new possibilities across neurology, psychiatry, and continuous patient monitoring [6]. Altogether, these advancements point to a future where AI supports both visual and signal-based insights, forming the backbone of smarter clinical decision-making tools.
Clinical decision support systems (CDSS) have become a vital part of current healthcare settings, offering insights drawn from electronic health records (EHRs) and real-time monitoring tools. Yet, many of the traditional AI models used in these systems fall short when it comes to flexibility, transparency, and oversight key qualities, especially critical in high-risk settings like intensive care units (ICUs). To address these limitations, we introduce a modular multi-agent system (MAS) designed to reflect how clinical teams make decisions, with a built-in emphasis on ethical AI to uphold both explainability and accountability.
Building on progress in large language model (LLM) agent-based frameworks, our system breaks down the decision-making pipeline into focused, collaborative agents. Each agent is responsible for a different aspect of ICU assessment: from interpreting lab results and tracking vital signs to making context-sensitive judgments based on a patient’s history or co-existing conditions. These individual agents pass their findings to an integration agent that brings everything together, enabling more comprehensive predictions, examining its transparency, and cross-validating its outcomes. This system simulates how doctors gather evidence from various sources, weigh context, and form a unified clinical picture.
By structuring the system around modular agents and grounding it in ethical oversight, we improve not just how interpretable the model is, but also how it upholds accountability throughout the clinical decision-making process. To test the framework, we used the eICU Collaborative Research Database [7], showing that our method can deliver well-organized predictions, shed light on key prognostic indicators, and build greater trust in AI-supported medical judgments.
@@ II-A Applications of Clinical Decision Support Systems in Intensive Care Settings
CDSS have come a long way, especially in ICU environments where every second counts. Earlier systems typically leaned on rule-based logic or statistical methods to generate recommendations [8, 9]. More recent developments have looked to clinical practice guidelines (CPGs) as a way to enrich LLMs, boosting their ability to offer context-aware treatment advice. Research suggests that LLMs enhanced with CPGs outperform traditional models in delivering more accurate clinical suggestions [8].
Meanwhile, MAS to CDSS have been gaining popularity. Researchers developed a particularly interesting framework that combines case-based reasoning with a MAS. This system uses different agents to manage how users interact with it, execute tasks, and apply medical knowledge[9]. What makes this approach valuable is how it merges MAS with case-based reasoning, allowing the system to learn more efficiently and better adapt to each patient’s unique situation.
@@
eICUs represent a significant influence on critical care medicine, harnessing telemedicine to address the shortage of on-site specialists for high-risk patients[11]. This innovative method enables continuous expert monitoring and intervention without the limitation of physical distance. For example, Philips’ eICU system uses the eCareManager platform to virtually bring ICU specialists to the patient’s bedside. By connecting hospital networks and providing real-time clinical feedback, eICUs effectively bridge the gap between remote experts and immediate patient needs [12].
The implementation of eICU has been notable, as evidenced by the experience at Baptist Health South Florida. After introducing their eICU model, the institution saw a significant 23% decrease in ICU mortality rates and up to a 25% reduction in average length of stay (LOS)[12]. These improvements really show how telemedicine is changing critical care for the better. The eICU approach makes care better in several important ways - doctors can watch patients around the clock, catch problems earlier, make better use of limited specialists, follow consistent treatment plans, and work more closely with the nurses and doctors at the bedside. Hospitals also receive financial benefits since patients are getting out of the ICU faster. In addition, all the data these systems collect is valuable for research, which can help to keep making ICU care better over time.
@@ II-C LLM-Based Agents in Healthcare
LLMs have recently become more common in healthcare. These models now assist in multiple areas, including virtual assistants, individualized health education, symptom checking, and mental health support tools [13]. By improving patient interactions and simplifying administrative tasks, LLM-based systems are beginning to influence how healthcare is provided.
One example is MDAgents, a MAS using LLMs to manage complex medical decisions [14]. Its design replicates the teamwork observed in actual healthcare environments, enabling effective communication among its agents. Testing has shown that MDAgents performs better than earlier models in various evaluations.
A recent review explored the use of LLM-based agents in medical contexts [15]. The review covered technical foundations, practical applications, and existing challenges. It emphasized components such as planning techniques, reasoning strategies, integration of external tools, and agent architecture. These systems are now employed for tasks like CDSS, automatic patient documentation, simulation training, and workflow optimization.
However, MedAgentBench, offers a benchmark for evaluating LLMs as medical agents, featuring 300 clinically-derived tasks across 10 categories and 100 realistic patient profiles. The results indicate that current LLMs still struggle with complex tasks, leading to the need for optimization prior to used in autonomous healthcare applications[16].
Therefore, LLM-based agents have been further developed into MAS, where multiple agents interact in a collaborative manner. This change allows for systems that are more organized and flexible, offering new ways to manage challenging healthcare situations, such as emergency response coordination and personalized patient treatment.
@@ II-D Multi-Agent Systems in Healthcare
MAS are gaining attention as a promising way to tackle complex challenges in healthcare. One example applies MAS to pre-hospital emergency response, where agents—such as EMS dispatch centers, ambulances, traffic nodes, and medical providers—collaborate within a distributed decision-making setup [17].
The idea of multi-stage AI agents builds on this by organizing intelligent agents into layers, each handling different parts of perception and reasoning. Many of these layers are now powered by LLMs, allowing for more structured and scalable workflows [18]. This layered setup has shown promise in areas like personalized care and remote health services. Furthermore, the LLM-medical-agent framework, for instance, demonstrates how MAS can be applied to modular analysis of healthcare data in practical settings [19].
@@ II-E Ethical Governance in Healthcare AI
Ethical rules applied in healthcare AI is addressed through Explainable AI (XAI), which showcases the decision-making processes of complex algorithms. ”Healthcare AI Datasheets” framework, which documents potential biases through demographic data[20]. These complementary approaches not only enhance transparency by demystifying the ”black-box” problem but also actively work toward equitable healthcare outcomes by identifying and addressing the sampling and complexity biases that have historically perpetuated healthcare disparities across diverse populations[21].
Growing concerns around the safety of LLM-based agents have prompted the development of MAS [22, 23], which embed ethical advisor and policy guardrails to ensure compliance with safety and privacy standards—an especially important safeguard in clinical environments [24].
@@
Thus, the World Health Organization (WHO) has outlined ethical guidelines for AI in healthcare, highlighting principles such as human autonomy, well-being, and system transparency [26]. Especially in high-risk areas like intensive care units, solving these governance challenges is key to the responsible deployment of AI [27].
@@ II-F Motivation and Research Gap
While AI-powered CDSS have been promisingly adopted, there are still needs to be considered for real-world ICU practices. Many solutions fall short in modularity due to lack of transparency, or aren’t built with an inter-agent communication system to reflect the dynamic, interdisciplinary nature of intensive care.
Currently, most existing approaches tend to focus on isolated tasks—like interpreting lab results, monitoring vital signs, or reasoning based on medical history—but few bring these components together into a unified or dynamic system that reflects how clinicians actually work as a team. The essential needs draw our attention to propose a novel agent-based system that breaks down the clinical reasoning process.
Thus, we build the system with specialized, collaborative agents—each designed to handle a distinct aspect of care while maintaining accountability and interpretability throughout. By applying ethical AI principles at every stage of the pipeline and validating our design using the eICU database, this study aims to bridge both the technical and ethical gaps in deploying trustworthy AI for high-stakes decision-making in critical care.
Figure 1: Illustration of the MAS Design. The system consists of a set of specialized agents, each responsible for processing a specific type of clinical data. The Context Analysis Agent handles unstructured inputs like clinical notes, while the Vitals Analysis Agent focuses on real-time physiological signals, and the Lab Analysis Agent interprets laboratory test results. These distinct streams of information are brought together by the Integration Agent, which fuses multimodal features into a unified representation. Based on this, the Prediction Agent carries out key forecasting tasks—such as predicting ICU mortality or estimating LOS. To support interpretability, the Transparency Agent generates human-readable, traceable explanations of model outputs. Finally, the Validation Agent oversees performance assessment by comparing predictions against ground truth data.
@@ III-A Dataset and Preprocessing
This study used the eICU Collaborative Research Database v2.0 [10], which compiles anonymized ICU records from over 200,000 patient admissions in different hospitals across the U.S. The database contains two types of data: 1) structured details (e.g., vital signs and lab results) and 2) unstructured clinical notes contributed by nurses and physicians, giving us a overall view of patient care.
Here we used several key eICU files: patient.csv, lab.csv, vitalPeriodic.csv, note.csv, and medication.csv, along with APACHE-related data files [28] (apacheApsVar.csv and apachePatientResult.csv). To align each patient’s information, records were grouped according to patient-unit-stayid. If any essential data were missing—such as vital signs, lab values, or clinical notes—we removed those entries to maintain reliability.
We filled in the data gaps, ordered events based on their timestamps, and shortened lengthy text fields to meet the language model guidelines. Then, we sampled 150 patients for the study: 76 mortality patients and 74 survived patients. This balanced sample, provided a controlled dataset for our comparative analysis.
Then, we retrieved specific features from each patient’s record set. We collected the ten most recent vital sign readings to reveal each patient’s current physiological state. We also selected the latest distinct lab biomarkers deemed clinically relevant. When dealing with unstructured clinical documentation, we included up to three notes for every patient, focusing primarily on entries written by physicians and nurses. In addition, our analysis tracked the top 20 medications and treatments, identifying them by frequency or uniqueness within the overall dataset. Finally, we incorporated APACHE scores and predictions as reference points to aid in validating and evaluating our modeling outcomes.
III-B Multi-Agent System Design
To emulate real-world ICU decision-making, we implemented a modular MAS consisting of six discrete agents, each responsible for a semantically distinct task. The system is shown in Figure 1.
@@ III-B Multi-Agent System Design
  • Integration Agent: Aggregates all the results from the above agent into a comprehensive, system-by-system clinical assessment. It prioritizes ICU risk factors related to mortality and length of stay (LOS).
  • Prediction Agent: Generates structured outcome predictions (mortality probability and ICU LOS) using integrated findings and APACHE variables. Results follow a strict template for automated parsing.
  • Transparency Agent: Analyzes clinical prediction outputs to meet ethical standards and are explainable to various stakeholders. This score includes calculating explainability, interpretability, and traceability scores based on specific steps for each dimension.
@@
To ensure information transfer between agents, we implemented a shared memory architecture that allows any agent to access inputs and outputs from previous pipeline stages. This approach reduces the risk of information loss between modules while maintaining the semantic separation of responsibilities. While this shared memory design has proven effective at improving predictive performance by ensuring that no critical data leakage during the analysis process, we recognize that the MAS introduces additional complexity that can impact transparency. Our current implementation focuses on performance optimization, with ongoing work to enhance the clarity across the agent communication pathways.
@@ III-D System Execution and DAG Orchestration
  • prediction feeds into transparency.
  • transparency feeds into validation.
@@
Figure 2: Performance comparison between the MAS and SAS across two evaluation metrics: Mortality Prediction Accuracy and Length of Stay (LOS) Mean Error. Each model was executed 8 times, and the box plots represent the distribution of results across these runs. The MAS demonstrates slightly higher mean mortality accuracy and consistently lower LOS errors than the SAS.
@@ III-F Ethical AI and Explainability
Here we incorporate the explainability,interpretability, and traceability aspects within our agent system:
  • Explainability: Measures how the system detects and explains the factors influencing predictions. This includes elements such as the patient’s age, intubation status, creatinine levels, and other clinical parameters.
  • Interpretability: Evaluates how the reasoning process is communicated among various stakeholders. This step can allow the decision-making logic is understandable to clinicians, patients, and administrators.
  • Traceability: Assesses the trace process starting from input data, as well as influencing factors back to their sources. This builds a solid trail of evidence backing up the system’s findings, making it easier to trust and use in real-world scenarios.